OpenStudy (anonymous):

If f(x) = −4x + 1, find f(−5). A.20 B.-19 C.21 D.-8

OpenStudy (anonymous):

I Got C but is that correct?

jimthompson5910 (jim_thompson5910):

f(-5) means "find f(x) when x = -5" f(x) = −4x + 1 f(-5) = −4(-5) + 1 ... replace all copies of 'x' with '-5' (since x = -5) f(-5) = 20 + 1 f(-5) = 21 so you are 100% correct

OpenStudy (anonymous):

basically its pretty much just saying fill in for x which is (-5) -4(-5)+1 20+1 = 21 C is the correct answer.

OpenStudy (anonymous):

@jim_thompson5910 & @genson0 thanks

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

You're welcome :) any time.